Apr 21, 2016 Calorimeter

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11:00 a.m.

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/907185247.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Action Items

  1. Modify covariance matrix for FCAL and BCAL to correspond to the current resolution of the detectors (Matt). Check pull distributions with new clusterizer.


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book
  3. BCAL Action Items and Working Tasks

Tentative Agenda

  1. Announcements
  2. Collaboration Meeting
  3. Action Items
  4. Run Preparations
    1. Online Monitoring Updates
    2. FCAL Update
    3. BCAL Update
  5. Reconstruction and Simulation
  6. Calibration
  7. Any other business


Attendees: Elton, Simon, Adesh, Mark (JLab); Curtis, Mike, Will (CMU); Ahmed, Andrei, Zisis, Tegan (UofR); Matt (IU); Christina (Athens); Sean (NW).

  1. Announcements
    • Magnetic field is being ramped to 1350 A this morning.
  2. Collaboration Meeting
    • Talks by Will (Bcal clusterizer, calibration); Adesh (Fcal); Ahmed/Andrei/Zisis (Regina updates); George (TDC veff, offsets)
  3. Action Items
    • Cal covariance matrix - reevaluate pull distributions with new clusterizer and calibrations
  4. Run Preparations
    1. Online Monitoring Updates
    2. FCAL Update
      1. Bases
        • Six bases replaced yesterday (2 HV, 4 communication problems).
        • The four with communication did not have their LED light on. Problem required recycling the power.
        • After resetting power 5 other bases did not come back.
        • At the moment, one cannot recycle the power supply remotely (only the ANA gates). A remotely controllable power strip will be installed this summer.
      2. Test setup
        • Setup at JLab is fully functional
        • Found current leakage problems and they are now fixed.
        • Base monitoring is ongoing. Only about 30% of failed based have shown problems in the lab
        • A new communication issue has arisen in recent failures
        • Curtis: How many bases to date have been replaced? Adesh: About 300.
        • Zisis: How many replacements have failed again? Adesh: Very few repeat failures. But some bases fixed for one reason have failed for other problems.
        • Matt: Note that many bases have been replaced for intermittent problems, which confuse the repair history. They are tricky to understand and hard to reproduce.
      3. HV settings / calibration
        • Original HV settings till Mar 23. Initial HV change (based on faulty evaluation) Mar 23-April 9. Latest HV settings since April 9.
        • Shortest period is Mar 23-Apr 9, but there are about 10M pi0 events, which should be sufficient for a good calibration
        • Mark Ito has set Adesh up to be able to run on the FCAL account (as well as his own).
        • Plan to have calibrations for each period by collaboration meeting.
        • Looking at pedestal stability and timing offsets.
      4. Pulser data
        • Latest production runs include the FCAL pulser, recording events at 1 Hz.
        • Using this data to check response to magnetic fields, including current increase of current from 1200 to 1350 A this morning.
    3. BCAL Update
  5. Reconstruction and Simulation
    1. mcsmear updates - BCAL, FCAL
      • Bcal comparison will wait for latest results with new clusterizer
      • Fcal data gives better resolution than the MC. Checked for threshold effects. It appears the MC has higher thresholds than the data (leading to worse resolution). Recheck with new HV values
      • Matt: Notes that the MC includes shower fluctuations, but the effect of statistics are added by hand in mcsmear, so this may be a reason for discrepancies.
    2. Bcal (Mark): Will has committed the first attempt to fix clusterizer and is much improved. The new version is included in the new software launch. Further refinements may be made
      • Mark has been looking at timing but is distracted as RC for this week.
    3. BCAL LED pulser
      • Ahmed is looking at near/far ratios and checking scripts
      • Elton: Has he looked at any of the production data? No, for the moment they are testing software on LED calibration runs.
      • Adesh: Can use a flag to tag only BCAL data. Mark: For these studies select events using front panel trigger bit.
    4. Installing sim-recon on Canadian cluster
      • Tegan has been working on installing GlueX software on Canadian cluster
      • Ran into some issues, but they are working through them.
    5. Andrei: No updates on time-walk corrections
    6. Sean: Is there any new effort to find shower efficiencies?
      • Mark: Plugging to look at single-channel efficiencies with charged particles is running during the launches.
      • Some ideas for looking at shower-level efficiencies include selecting pi0->g e+e-, which Will started working on but was diverted to the clusterizer. He will get back to this.
  6. Calibration
  7. Any other business
    • Zisis will run the next meeting.